# Towards an exact electronic quantum many-body treatment of Kondo correlation in magnetic impurities

### v1

@@ Abstract
The Kondo effect is a prototypical quantum phenomenon arising from the interaction between localized electrons in a magnetic impurity and itinerant electrons in a metallic host. Although it has served as the testing ground for quantum many-body methods for decades, the precise description of Kondo physics with material specificity remains challenging. Here, we present a systematic ab initio approach to converge towards an exact zero-temperature electronic treatment of Kondo correlations. Across a series of $3d$ transition metals, we extract Kondo temperatures matching the subtle experimental trends, with an accuracy far exceeding that of standard models. We further obtain microscopic insight into the origin of these trends. More broadly, we demonstrate the possibility to start from fully ab initio many-body simulations and push towards the realm of converged predictions.
@@ Introduction
The challenge of the Kondo problem stems from the simultaneous presence of strong electron correlation leading to local singlet formation, and impurity-metal hybridization, which requires considering the thermodynamic limit. In addition, the Kondo energy scale is very small (typically 1 to 500 K). In low-energy models, one chooses a (small) set of impurity correlated orbitals (e.g., a few $d$ or $f$ orbitals), whose interactions are formally described by a downfolded effective Hamiltonian. Uncertainties, which are large on the Kondo scale, then arise from the choice and construction of these orbitals [Karp2021a], the derivation of the effective interactions (and approximations to treat their complicated frequency dependence [Aryasetiawan2004b]), and the use of double-counting corrections to remove density functional theory (DFT) contributions [Wang2012, Muechler2022b], in lieu of a fully many-body treatment of the local interactions. As a result, although much qualitative progress has been made in describing Kondo physics in different realistic settings [Jacob2009, Surer2012a, Gardonio2013, Dang2016a, Valli2020], the quantitative simulation of Kondo trends across different magnetic impurities, geometries, or even different calculations, without reference to experimental data, is challenging [Surer2012a, Dang2016a, Gardonio2013].
We will instead attempt to solve the ab initio many-electron Schrödinger equation for the Kondo problem without first deriving a low-energy model. This offers the advantage that it is often easier to quantify (and thus converge) errors in the approximate solution of a problem, than the error associated with deriving a model. We build on our work on full cell embedding in the context of dynamical mean-field theories [Zhu2020, Cui2020, Zhu2021c, Cui2022, Cui23-cuprate-doping] to construct a systematically improvable representation of the impurity in its metallic host. Solving the many-body problem in this representation, as we increase the number of orbitals in the parent basis, we eventually obtain the exact non-relativistic, zero-temperature, description of the pure electronic problem. In small molecules [larsson2022chromium], as well as in simple materials (such as organic crystals [yang2014ab]), it has been established that for the quantities of interest, related strategies can reach an accuracy rivalling, or even exceeding, that of experimental measurements. We thus ask, how far can we go with a similar philosophy for a realistic correlated quantum materials problem?
@@ Results
More specifically, we study the series of seven $3d$ transition metal impurities (Ti, V, Cr, Mn, Fe, Co, Ni) in bulk Cu, with associated Kondo temperatures that span three orders of magnitude. We started from a Gaussian atomic orbital representation of all atoms: a split-valence double-$\zeta$ polarization basis (def2-SVP) [Weigend2005] for the impurity atoms and a slightly smaller (def2-SV(P)) basis for Cu (which omits the $4f$ basis function). The def2-SVP basis for the impurity atoms contains the $1s2s2p3s3p3d4s4p4d4f5s$ shells (denoted a $5s3p2d1f$ basis, from the shell-count) and thus goes significantly beyond the $3d$ shell considered in a model Hamiltonian treatment. To test the convergence of the results, for a subset of the calculations we also used a larger cc-pVTZ basis [Balabanov2005a] on the impurity atom, corresponding to a $7s6p4d2f1g$ basis. This further improves the electronic treatment of the impurity. We will refer to these as the parent bases and below we demonstrate convergence towards the exact solutions within the parent bases. We estimate the remaining error of the parent basis by the difference between the def2-SVP and cc-pVTZ results.
Starting from DFT-optimized XCu63 structures (X = impurity), we performed periodic DFT calculations in the Gaussian atomic orbital bases with the PBE functional [Perdew96PBE, Sun2020b]. The Gaussian basis functions were then transformed into an orthogonal basis of intrinsic atomic orbitals plus projected atomic orbitals (IAO+PAO) [Knizia2013d]. The impurity IAOs and PAOs are visualized in Fig. 1a. The higher shell orbitals in this picture extend away from the impurity atom onto the neighbours, and because of the close packing of the atoms, start to capture important electronic degrees of freedom of the atoms neighbouring the impurity. For example, the Löwdin population of the spatial extent of the $4s$ orbital of an Fe impurity on the neighbouring Cu atoms is close to $66\%$ in FeCu63. The large basis sets may thus be viewed as forming an “impurity-centered” basis expansion, similar to well-studied local correlation treatments in quantum chemistry.
@@
where the impurity sum extends over all IAOs and PAOs in the impurity basis, $\{a^{(\dagger)}_{i}\}$ are creation/annihilation operators for impurity states, and $\{c^{(\dagger)}_{p}\}$ are creation/annihilation operators for bath states. The impurity Coulomb interaction matrix $(ij|kl)$ is taken as the bare Coulomb interaction in the IAO+PAO basis, while the impurity one-particle interaction $\tilde{F}_{ij}$ in Eq. 1 is defined as the Hartree-Fock effective Hamiltonian with the local mean-field potential subtracted (this subtraction is diagrammatically exact, so there is no double counting). The impurity-metal hybridization function $\Delta(\omega)$ was obtained at the PBE level. Since we used a Hamiltonian-based impurity solver, we discretized the $3d4s$ valence block of the hybridization function on a non-uniform grid along the real frequency axis [DeVega2015c]
@@
Figure 1: Illustration and validation of computational strategy. (a) All-orbital quantum embedding strategy, where general Coulomb interactions between all valence and high-virtual orbitals of the impurity atom are treated at the quantum many-body level. The percentage of orbital extension onto neighbouring Cu atoms is estimated through the Löwdin population analysis. (b) Experimentally measured Kondo temperature trend ($T_{\mathrm{K}}$ averaged over available experimental values [Daybell1968]) for $3d$ transition metal impurities in bulk Cu, plotted versus the nominal spin ($S$) of isolated impurity atoms. (c) Simulated Kondo resonance and self-energies of the $t_{2g}$ and $e_{g}$ orbitals of the Fe impurity with an active space of (22e, 22o) in the def2-SVP basis. (d) Estimated convergence of the quasiparticle renormalization weights and orbital occupancies of the $3d$ orbitals of Ti, V, and Fe impurities as the DMRG active space increases in the def2-SVP basis. (e) Comparison of quasiparticle renormalization weights and orbital occupancies of $3d$ orbitals of the Fe impurity calculated in the def2-SVP basis versus the cc-pVTZ basis.
To compute and converge the impurity Green’s function, we constructed a series of smaller orbital spaces (active spaces) that are subspaces of the full orbital space of the embedding problem. We then monitored the convergence of the ground-state and dynamical quantities as a function of the active space size. The active spaces were defined to span natural orbitals (eigenvectors of the one-particle density matrix) of an (approximate) ground state of the impurity problem, taking natural orbitals with the largest orbital entropy (a procedure which has been shown to yield highly compact orbital subspaces in model impurity problems [Zgid2012]). We used active spaces of up to (36e, 52o) for the ground-state problem in the impurity def2-SVP basis and (46e, 76o) in the impurity cc-pVTZ basis, and up to (26e, 36o) for the dynamical properties. We computed ground states and associated properties (e.g. density matrices and correlation functions) using ab initio DMRG [Chan2011a, Zhai2021a, Zhai2023a]. For dynamical properties, we used a new zero-temperature ab initio dynamical DMRG (DDMRG) solver to compute the impurity self-energy on the real frequency axis [Ronca2017b, Zhai2023a]. (Benchmarks of the DDMRG solver on the single-impurity Anderson model are shown in Supplementary Note 2).
@@
We first consider the Fe impurity to illustrate general features of the results, then we discuss numerical convergence. In Fig. 1c we show the orbital-resolved excitation spectra and self-energies of the Fe $3d$ orbitals. As expected, we see sharp Kondo resonance peaks around the Fermi level for both the $t_{2g}$ and $e_{g}$ orbitals, with broad Hubbard shoulder peaks at higher energies. We find that the Fe-$e_{g}$ Kondo peak is narrower and smaller than the Fe-$t_{2g}$ peak from the stronger QP renormalization, which can also be seen from the real part of the real-axis self-energies shown in Fig. 1c. The imaginary part of the real-axis self-energies of both the $t_{2g}$ and $e_{g}$ orbitals approaches zero at the Fermi level (Fig. S5), i.e., $\mathrm{Im}\Sigma(T=0,\omega)\rightarrow 0$ as $\omega\rightarrow 0$. The splitting of the Hubbard peaks is roughly $2.5\sim 3$ eV (in the range of screened interaction for Fe $3d$ orbitals [Nakamura2008a]) and is slightly smaller for $t_{2g}$ orbitals than for $e_{g}$ orbitals.
@@
We then examine the convergence of impurity observables towards the parent basis limit. In Fig. 1d, we show the orbital-resolved QP renormalization weight $Z$ (calculated on real axis as $\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}$) and the orbital occupancy $n$ of the $3d$ orbitals, for the Ti, V, and Fe impurities, as a function of active space size with a parent def2-SVP basis. Additional convergence data is in the Supplementary Notes 3 and 4, including the convergence with respect to the DMRG bond dimension (for which the uncertainty is substantially smaller than the uncertainty from the active space for all cases other than Mn). We observe that the QP weight is clearly more challenging to converge. Examining the full series of elements (Supplementary Notes 3 and 4) we conservatively estimate that, with the exception of Mn, $n$ is converged to at least $\sim 0.04$ and $Z$ to at least $\sim 0.05$ of the exact parent basis result. The latter exception arises because it was difficult to fully converge the DMRG calculations in the largest active spaces for Mn.
@@
Beyond these errors, the primary source of deviation from an exact simulation comes from the insufficiency of the parent basis itself. To estimate the error from the parent basis, in Fig. 1e we show the change in $Z$ and $n$ as we significantly increase the parent basis from def2-SVP to cc-pVTZ, for the Fe impurity. We see that both quantities change by less than $\sim 0.03$. Standard quantum chemistry arguments suggest that most quantities converge like the inverse cube of the cardinal number of the basis [helgaker1997basis] (2 in the case of def2-SVP, 3 for cc-pVTZ), which indicates that $Z$ and $n$ are well converged with respect to the representation of the electronic degrees of freedom of the Fe impurity and its close neighbours. Only the contribution from fluctuations involving long-range Coulomb interactions beyond the closest neighbours remains outside of our treatment and this error analysis.
Quantitative Kondo temperature trends and orbital-specific physics.
Having examined the convergence of our numerical results, we now study the trends in Kondo physics across the series of 7 elements. For $3d$ magnetic impurities in bulk Cu, experiments observe an exponential decay in Kondo temperatures moving from nominally low-spin (e.g., Ti) to high-spin (e.g., Mn) impurities [Daybell1968]. This suppression of the Kondo temperature is understood to result from the reduction of the effective exchange coupling $J_{\mathrm{eff}}$, due to Hund’s rule coupling [Georges2013]. Fixing the metallic host, and assuming all other effects are the same for different impurities, this relation can be simplified to
@@
where $S$ is the spin of the magnetic atom. The estimated experimental Kondo temperatures for the $3d$ magnetic impurities, along with the nominal spin $S$ of the atom is shown in Fig. 1b. Note that the experimental estimates of $T_{\mathrm{K}}$ involve significant interpretation (for a full discussion see [Daybell1968]) and thus a range of characteristic temperatures are obtained depending on the type of measurement and method to extract $T_{\mathrm{K}}$. Nonetheless, Eq. 3 clearly misses quantitative differences between certain magnetic impurities. For example, it does not explain why $T_{\mathrm{K}}$ of Cr (3$\sim$10 K) is lower than $T_{\mathrm{K}}$ of Fe (30 K), if we consider both Cr and Fe atoms to have $S=2$.
To obtain a computational perspective using our methodology, in Fig. 2a we first show the density of states of the Cr, Mn, and Fe impurities. The $t_{2g}$ and $e_{g}$ orbitals of all three impurities yield narrow Kondo resonance peaks around the Fermi level. Since the Kondo temperature $T_{\mathrm{K}}$ is proportional to the width of the Kondo resonance peak, Fig. 2a suggests that the Kondo temperatures of the $t_{2g}$ and $e_{g}$ orbitals follow different trends: $T_{\mathrm{K}}(\mathrm{Cr})<T_{\mathrm{K}}(\mathrm{Mn})<T_{\mathrm{K}}(\mathrm{Fe})$ for the $t_{2g}$ orbitals and $T_{\mathrm{K}}(\mathrm{Mn})<T_{\mathrm{K}}(\mathrm{Fe})\approx T_{\mathrm{K}}(\mathrm{Cr})$ for the $e_{g}$ orbitals. This is corroborated by the behaviour of the real-frequency self-energies (Fig. S5), and illustrates that the trend across the elements is intertwined with orbital-specific physics.
Figure 2: Trends in impurity DOS, quasiparticle renormalization, spin moment, and Kondo temperature. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) and Green’s function active space of (22e, 22o). (a) Comparison of DOS of $3d$ orbitals of Cr, Mn, and Fe impurities. (b) Top: Orbital-resolved quasiparticle renormalization weights $Z$. Bottom: Local spin moments of impurity atoms compared to nominal spins of isolated atoms and experimental values. (c) Kondo temperatures from all-orbital DMFT predictions compared to experimental values [Daybell1968]. The experimental trend lines are provided as a visual aid. Symbols for each type of measurements are: $\rho$: resistivity, $\chi$: susceptibility, $C$: specific heat, $S$: thermoelectric power.
For a more quantitative estimate, we extracted Kondo temperatures for the $t_{2g}$ and $e_{g}$ orbitals for all seven $3d$ magnetic impurities using the relation from Hewson’s renormalized perturbation theory of the Anderson impurity model [Hewson1993]
$T_{\mathrm{K}}=-\frac{\pi}{4}Z\cdot\mathrm{Im}\Delta(0).$ (4)
Here $\Delta(0)$ is the hybridization function at the Fermi level and $Z$ is the QP renormalization weight on the real axis (Fig. S4 and Table S4). $\Delta(0)$ only varies slightly across the series (from $-0.59$ eV (Ti) to $-0.43$ eV (Fe) for the $t_{2g}$ orbital and from $-0.64$ eV (Ti) to $-0.38$ eV (Ni) for the $e_{g}$ orbital). On the other hand, we see a dramatic reduction of the QP renormalization weight $Z$ moving from low-spin to high-spin impurities for both the $t_{2g}$ and $e_{g}$ orbitals, as seen in Fig. 2b, and the trend is orbital-dependent. For the $t_{2g}$ orbitals, Cr has the lowest QP weight, at $Z=0.017$, and the trend for QP weight follows: Cr $<$ Mn $<$ Fe $<$ Co $\approx$ V $<$ Ni $<$ Ti. For the $e_{g}$ orbitals, Mn has the strongest QP renormalization with $Z=0.006$ and the trend for QP weight follows: Mn $<$ Fe $<$ Cr $<$ Co $<$ V $<$ Ni $<$ Ti.
From Eq. 4, we obtain $T_{\mathrm{K}}(t_{2g})=73\leavevmode\nobreak\ \mathrm{K}$ and $T_{\mathrm{K}}(e_{g})=167\leavevmode\nobreak\ \mathrm{K}$ for Cr, and $T_{\mathrm{K}}(t_{2g})=565\leavevmode\nobreak\ \mathrm{K}$ and $T_{\mathrm{K}}(e_{g})=101\leavevmode\nobreak\ \mathrm{K}$ for Fe. The lower of the two Kondo temperatures is the relevant one for the measurements (see e.g., Ref. [daybell1967observation] for resistivity), thus we obtain characteristic Kondo temperatures of Cr (73 K) and Fe (101 K). This indeed reproduces the subtle experimental trend, which we see is driven by electron correlation in different orbitals in Cr ($t_{2g}$) and Fe ($e_{g}$) respectively.
@@
More insights can be obtained from the computed local ground-state properties of the impurity, including the $t_{2g}$ and $e_{g}$ orbital occupancies across the impurity elements (Table S5) and the effective spin (extracted from $\langle S^{2}\rangle$) in Fig. 2b. Charge transfer takes place to the impurity atom in all systems, giving negatively charged $3d$ shells compared to the isolated impurities. For instance, we find Fe in bulk Cu to be on average in a $d^{7}$ occupancy ($n_{\mathrm{Fe}}(3d)=6.92$), in excellent agreement with recent experimental estimates [Joly2017]. The strong QP renormalization in Fig. 2b tracks the partially-filled nature of the impurity $3d$ orbitals. For example, we obtain $n(t_{2g})=0.99$ in Cr and $n(e_{g})=1.03$ in Mn (i.e., very close to half filling), which correlates well with their very small QP weights.

### v3

@@ Abstract
The Kondo effect is a prototypical quantum phenomenon arising from the interaction between localized electrons in a magnetic impurity and itinerant electrons in a metallic host. Although it has served as the testing ground for quantum many-body methods for decades, the precise description of Kondo physics with material specificity remains challenging. Here, we present a systematic ab initio approach to converge towards an exact zero-temperature electronic treatment of Kondo correlations. Across a series of $3d$ transition metals, we extract Kondo temperatures matching the subtle experimental trends, with an accuracy exceeding that of standard models. We further obtain microscopic insight into the origin of these trends. More broadly, we demonstrate the possibility to start from fully ab initio many-body simulations and push towards the realm of converged predictions.
@@ Introduction
The challenge of the Kondo problem stems from the simultaneous presence of strong electron correlation leading to local singlet formation, and impurity-metal hybridization, which requires considering the thermodynamic limit. In addition, the Kondo energy scale is very small (typically 1 to 500 K). In low-energy models, one chooses a (small) set of impurity correlated orbitals (e.g., a few $d$ or $f$ orbitals), whose interactions are formally described by a downfolded effective Hamiltonian. Uncertainties, which are large on the Kondo scale, then arise from the choice and construction of these orbitals [Karp2021a], the derivation of the effective interactions (and approximations to treat their complicated frequency dependence [Aryasetiawan2004b]), and the use of double-counting corrections to remove density functional theory (DFT) contributions [Wang2012, Muechler2022b], in lieu of a fully many-body treatment of the local interactions. As a result, although much qualitative progress has been made in describing Kondo physics in different realistic settings [Jacob2009, Surer2012a, Gardonio2013, Dang2016a, Valli2020], the accurate simulation of Kondo trends across different magnetic impurities, geometries, or even different calculations, without reference to experimental data, is challenging [Surer2012a, Dang2016a, Gardonio2013].
We will instead attempt to solve the ab initio many-electron Schrödinger equation for the Kondo problem without first deriving a low-energy model. This offers the advantage that it is often easier to quantify (and thus converge) errors in the approximate solution of a problem, than the error associated with deriving a model. We build on our work on full cell embedding in the context of dynamical mean-field theories [Zhu2020, Cui2020, Zhu2021c, Cui2022, Cui23-cuprate-doping] to construct a systematically improvable representation of the impurity in its metallic host. Solving the many-body problem in this representation, as we increase the number of orbitals in the parent basis, we eventually obtain the exact non-relativistic description of the pure electronic problem. In small molecules [larsson2022chromium], as well as in simple materials (such as organic crystals [yang2014ab]), it has been established that for the quantities of interest, related strategies can reach an accuracy rivalling, or even exceeding, that of experimental measurements. We thus ask, how far can we go with a similar philosophy for a realistic correlated quantum materials problem?
@@ Results
More specifically, we study the series of seven $3d$ transition metal impurities (Ti, V, Cr, Mn, Fe, Co, Ni) in bulk Cu, with associated Kondo temperatures that span three orders of magnitude. We started from a Gaussian atomic orbital representation of all atoms: a split-valence double-$\zeta$ polarization basis (def2-SVP) [Weigend2005] for the impurity atoms and a slightly smaller (def2-SV(P)) basis for Cu (which omits the $4f$ basis shell). The def2-SVP basis for the impurity atoms contains the $1s2s2p3s3p3d4s4p4d4f5s$ shells (denoted a $5s3p2d1f$ basis, from the shell-count) and thus goes significantly beyond the $3d$ shell considered in a model Hamiltonian treatment. To test the convergence of the results, for a subset of the calculations we also used a larger cc-pVTZ basis [Balabanov2005a] on the impurity atom, corresponding to a $7s6p4d2f1g$ basis. This further improves the electronic treatment of the impurity. We will refer to these as the parent bases and below we demonstrate convergence towards the exact solutions within the parent bases. We estimate the remaining error of the parent basis by the difference between the def2-SVP and cc-pVTZ results.
Starting from DFT-optimized XCu63 structures (X = impurity), we performed periodic DFT calculations in the Gaussian atomic orbital bases with the PBE functional [Perdew96PBE, Sun2020b]. The Gaussian basis functions were then transformed into an orthogonal basis of intrinsic atomic orbitals plus projected atomic orbitals (IAO+PAO) [Knizia2013d]. The impurity IAOs and PAOs are visualized in Fig. 1a. The higher shell orbitals in this picture extend away from the impurity atom onto the neighbours, and because of the close packing of the atoms, start to capture important electronic degrees of freedom of the atoms neighbouring the impurity. For example, using the Löwdin population to measure the spatial extent of the $4s$ orbital of an Fe impurity, we find that it is close to $66\%$ on the neighbouring Cu atoms in FeCu63. The large basis sets may thus be viewed as forming an “impurity-centered” basis expansion, similar to well-studied local correlation treatments in quantum chemistry.
@@
where the impurity sum extends over all IAOs and PAOs in the impurity basis, $\{a^{(\dagger)}_{i}\}$ are creation/annihilation operators for impurity states, and $\{c^{(\dagger)}_{p}\}$ are creation/annihilation operators for bath states. The impurity Coulomb interaction matrix $(ij|kl)$ is taken as the bare Coulomb interaction in the IAO+PAO basis, while the impurity one-particle interaction $\tilde{F}_{ij}$ in Eq. 1 is defined as the Hartree-Fock effective Hamiltonian with the local mean-field potential subtracted (this subtraction is diagrammatically exact, so there is no double counting). All interactions are treated non-relativistically (spin-orbit coupling is beyond the scope of this work). The impurity-metal hybridization function $\Delta(\omega)$ was obtained at the PBE level. Since we used a Hamiltonian-based impurity solver, we discretized the $3d4s$ valence block of the hybridization function on a non-uniform grid along the real frequency axis [DeVega2015c]
@@
Figure 1: Illustration and validation of computational strategy. (a) All-orbital quantum embedding strategy, where general Coulomb interactions between all valence and high-virtual orbitals of the impurity atom are treated at the quantum many-body level. The percentage of orbital extension onto neighbouring Cu atoms is estimated through a Löwdin population analysis. (b) Experimentally measured Kondo temperatures for $3d$ transition metal impurities in bulk Cu. Each dot corresponds to a distinct measurement, while the shaded areas illustrate the range of Kondo temperatures from different observables and their associated measurement uncertainties (where applicable) [Daybell1968]. The $x$-axis shows the nominal spin ($S$) of the isolated impurity atoms. (c) Simulated Kondo resonance and self-energies of the $t_{2g}$ and $e_{g}$ orbitals of the Fe impurity with an active space of (22e, 22o) in the def2-SVP basis. (d) Estimated convergence of the quasiparticle renormalization weights and orbital occupancies of the $t_{2g}$ and $e_{g}$ orbitals, averaged within each manifold, of Ti, V, and Fe impurities as the DMRG active space increases in the def2-SVP basis. (e) Comparison of quasiparticle renormalization weights and orbital occupancies of $3d$ orbitals of the Fe impurity calculated in the def2-SVP basis versus the cc-pVTZ basis.
To compute and converge the impurity Green’s function, we constructed a series of smaller orbital spaces (active spaces) that are subspaces of the full orbital space of the embedding problem. We then monitored the convergence of the eigenstate and dynamical quantities as a function of the active space size. The active spaces were defined to span natural orbitals (eigenvectors of the one-particle density matrix) of an (approximate) ground state of the impurity problem, taking natural orbitals with the largest orbital entropy (a procedure which has been shown to yield highly compact orbital subspaces in model impurity problems [Zgid2012]). We used active spaces of up to (36e, 52o) for the ground-state problem in the impurity def2-SVP basis and (46e, 76o) in the impurity cc-pVTZ basis, and up to (26e, 36o) for the dynamical properties and the finite temperature simulations. We computed ground states and associated properties (e.g., density matrices and correlation functions), and as many as several hundred excited states, using ab initio DMRG [Chan2011a, Zhai2021a, Zhai2023a]. For dynamical properties, we used a new zero-temperature ab initio dynamical DMRG (DDMRG) solver to compute the impurity self-energy on the real frequency axis [Ronca2017b, Zhai2023a]. (Benchmarks of the DDMRG solver on the single-impurity Anderson model are shown in Supplementary Note 2).
@@
We first consider the Fe impurity to illustrate general features of the results. In Fig. 1c we show the orbital-resolved excitation spectra and self-energies of the Fe $3d$ orbitals. As expected, we see sharp Kondo resonance peaks, also known as Abrikosov-Suhl resonances [tsvelick1983exact], around the Fermi level for both the $t_{2g}$ and $e_{g}$ orbitals, with broad Hubbard shoulder peaks at higher energies. We find that the Fe-$e_{g}$ Kondo peak is narrower and smaller than the Fe-$t_{2g}$ peak from the stronger QP renormalization, which can also be seen from the real part of the real-axis self-energies shown in Fig. 1c. The imaginary part of the real-axis self-energies of both the $t_{2g}$ and $e_{g}$ orbitals approaches zero at the Fermi level (Fig. S5), i.e., $\mathrm{Im}\Sigma(T=0,\omega)\rightarrow 0$ as $\omega\rightarrow 0$. The splitting of the Hubbard peaks is roughly $2.5\sim 3$ eV (in the range of the screened interaction for Fe $3d$ orbitals [Nakamura2008a]) and is slightly smaller for the $t_{2g}$ orbitals than for the $e_{g}$ orbitals.
@@
We then examine the convergence of impurity observables towards the parent basis limit. In Fig. 1d, we show the orbital-resolved QP renormalization weight $Z$ (calculated on the real axis as $\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}$) and the orbital occupancy $n$ of the $3d$ orbitals, for the Ti, V, and Fe impurities, as a function of active space size with a parent def2-SVP basis. Additional convergence data are in the Supplementary Notes 3 and 4, including the convergence with respect to the DMRG bond dimension (for which the uncertainty is substantially smaller than the uncertainty from the active space for all cases other than Mn). We observe that the QP weight is clearly more challenging to converge. Examining the full series of elements (Supplementary Notes 3 and 4) we conservatively estimate that, with the exception of Mn, $n$ is converged to at least $\sim 0.04$ and $Z$ to at least $\sim 0.05$ of the exact parent basis result. The latter exception arises because it was difficult to fully converge the DMRG calculations in the largest active spaces for Mn.
@@
Beyond these errors, the primary source of deviation from an exact simulation comes from the insufficiency of the parent basis itself. To estimate the error from the parent basis, in Fig. 1e we show the change in $Z$ and $n$ as we increase the parent basis from def2-SVP to cc-pVTZ, for the Fe impurity. We see that both quantities change by less than $\sim 0.03$. Standard quantum chemistry arguments suggest that most quantities converge like the inverse cube of the cardinal number of the basis [helgaker1997basis] (2 in the case of def2-SVP, 3 for cc-pVTZ), which indicates that $Z$ and $n$ are well converged with respect to the representation of the electronic degrees of freedom of the Fe impurity and its close neighbours. Only the contribution from fluctuations involving long-range Coulomb interactions beyond the closest neighbours remains outside of our treatment and this error analysis.
Detailed Kondo temperature trends and mechanisms.
Having examined the convergence of our numerical results, we now study the trends in Kondo physics across the series of 7 elements. For $3d$ magnetic impurities in bulk Cu, experiments observe an exponential decay in Kondo temperatures moving from nominally low-spin (e.g., Ti) to high-spin (e.g., Mn) impurities [Daybell1968]. The estimated experimental Kondo temperatures for the $3d$ magnetic impurities, along with the nominal spins $S$ of the atom, are shown in Fig. 1b. Note that the experimental estimates of $T_{\mathrm{K}}$ involve significant interpretation (for a full discussion see Ref. [Daybell1968]) and thus a range of characteristic temperatures are obtained depending on the type of measurement and method to extract $T_{\mathrm{K}}$.
We estimate theoretical Kondo temperatures from Hewson’s renormalized perturbation theory of the Anderson impurity model [Hewson1993]
$T_{\mathrm{K}}=-\frac{\pi}{4}Z\cdot\mathrm{Im}\Delta(0).$ (3)
Here $\Delta(0)$ is the hybridization function at the Fermi level and $Z$ is the QP renormalization weight on the real axis (Fig. S4 and Table S6). Hewson’s formula gives separate estimates of $T_{\mathrm{K}}$ for the $e_{g}$ and $t_{2g}$ orbitals; the lower of the two Kondo temperatures is the relevant one for the comparison to measurements (see, e.g., Ref. [daybell1967observation] for resistivity).
Figure 2: Trends in impurity DOS, quasiparticle renormalization, spin moment, Kondo temperature, and orbital character of the lowest triplet excited state. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) and Green’s function active space of (22e, 22o). (a) Comparison of DOS of $3d$ orbitals of Cr, Mn, and Fe impurities. (b) Top: Orbital-resolved quasiparticle renormalization weights $Z$. Bottom: Local spin moments of impurity atoms compared to nominal spins of isolated atoms and experimental values. (c) Kondo temperatures from all-orbital simulations, estimated by zero-temperature Green’s function calculations using Hewson’s formula, compared to experimental values [Daybell1968]. The experimental trend lines are provided as a visual aid. Symbols for each type of measurement are: $\rho$: resistivity, $\chi$: susceptibility, $C$: specific heat, $S$: thermoelectric power. (d) Orbital-resolved spin-flip transition density matrix for the lowest triplet excited state $|\langle\Psi_{S=0}|c^{\dagger}_{i\downarrow}c_{j\uparrow}|\Psi_{S=1,S_{z}=1}\rangle|$ within the impurity $3d$ subspace.
Our predicted Kondo temperatures for all the studied elements are shown in Fig. 2c. We capture the trend in $T_{\mathrm{K}}$ across the full set of elements in good agreement with the experimental estimates, and reproduce subtle differences such as the relative ordering of the lowest Kondo temperatures $T_{\mathrm{K}}(\mathrm{Cr})$, $T_{\mathrm{K}}(\mathrm{Mn})$, $T_{\mathrm{K}}(\mathrm{Fe})$. With the exception of Cr and Mn, the range of our predicted $T_{\mathrm{K}}$ is within a factor of 2$\sim$3 of at least one of the experimental estimates of $T_{\mathrm{K}}$. As our calculations do not include long-range Coulomb fluctuations (beyond those captured by the basis extending to neighbouring atoms), this suggests that they do not contribute significantly at this level of accuracy. On the other hand, Mn, which has the lowest $T_{\mathrm{K}}$, is a system where we could not achieve numerical convergence of our results to the parent basis. This is the likely origin of the larger quantitative discrepancy for Mn.
Conventionally, the origin of the observed trend in $T_{\mathrm{K}}$ is ascribed to strong Hund’s coupling in the impurity spins. In particular, when we represent the system by a multi-spin Kondo model, in the strong coupling limit the impurity spins are locked, which reduces the space of fluctuations and the effective exchange coupling $J_{\mathrm{eff}}$ [okada1973singlet, Georges2013] (see Supplementary Note 15 for a brief summary). Fixing the metallic host, and assuming all other effects are the same for different impurities, this Hund’s driven relation can be simplified to
@@
where $S$ is the spin of the magnetic atom. The above theoretical form gives qualitative agreement with the experimental data, as shown by the trend in Fig. 1b. However, it also misses quantitative differences between certain magnetic impurities. For example, it does not explain why the $T_{\mathrm{K}}$ of Cr (3$\sim$10 K) is lower than the $T_{\mathrm{K}}$ of Fe (30 K), if we consider both Cr and Fe atoms to have nominal $S=2$. Given that our ab initio simulations faithfully reproduce the experimental trend in Kondo temperatures, we can now examine the Hund’s suppression mechanism within our calculations.
@@
We start with the computed local ground-state properties of the impurity, such as the $t_{2g}$ and $e_{g}$ orbital occupancies across the impurity elements (Table S7) and the effective spin (extracted from $\langle S^{2}\rangle$) in Fig. 2b. (Fig. 2a further illustrates the $t_{2g}$ and $e_{g}$ resolved DOS for Cr, Mn, Fe). Charge transfer takes place to the impurity atom in all systems, giving negatively charged $3d$ shells compared to the isolated impurities. For instance, we find Fe in bulk Cu to be on average in a $d^{7}$ occupancy ($n_{\mathrm{Fe}}(3d)=6.92$), in excellent agreement with recent experimental estimates [Joly2017]. The strong QP renormalization in Fig. 2b tracks the partially-filled nature of the impurity $3d$ orbitals. For example, we obtain $n(t_{2g})=0.99$ in Cr and $n(e_{g})=1.03$ in Mn (i.e., very close to half filling), which correlates well with their very small QP weights.